A department store sells eight ties and seven shirts for $246 how much Is one tie and one shirt
Answer by Edwin McCravy(20086)

There are 5 possibilities if both ties and shirts must sell
for a whole number of dollars 

If ties sell for $1 each and shirts sell for $34, then

8 ties sell for    $1×8 =   $8
7 shirts sell for $34×7 = $238
------------------------------
                          $246

And therefore one tie and one shirt costs $1 + $34 = $35


OR

If ties sell for $8 each and shirts sell for $26, then

8 ties sell for    $8×8 =  $64
7 shirts sell for $26×7 = $182
------------------------------
                          $246

And therefore one tie and one shirt costs $8 + $26 = $34

OR

If ties sell for $15 each and shirts sell for $18, then

8 ties sell for   $15×8 = $120
7 shirts sell for $18×7 = $126
------------------------------
                          $246

And therefore one tie and one shirt costs $15 + $18 = $33

OR

If ties sell for $22 each and shirts sell for $10, then

8 ties sell for   $22×8 = $176
7 shirts sell for $10×7 =  $70
------------------------------
                          $246

And therefore one tie and one shirt costs $22 + $10 = $32

OR

If ties sell for $29 each and shirts sell for $2, then

8 ties sell for   $29×8 = $232
7 shirts sell for  $2×7 = $ 14
------------------------------
                          $246

And therefore one tie and one shirt costs $29 + $2 = $31

So one tie and one shirt will cost between $31 and $35

Edwin